$SPAD/src/input d01bbf.input

The Axiom Team December 3, 2016

Abstract

Contents

2

1 License

— license —

--Copyright The Numerical Algorithms Group Limited 1994.

———-

— * —

\getchunk{license}
)clear all
showArrayValues true
showScalarValues true
a:SF:=0.0
b:SF:=1.0
itype:=1
n:=6
kind:=0
result:=d01bbf(a,b,itype,n,kind,-1)

———-

References^{[1] nothing}

4